# MASQ: Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion

### v1

@@ Abstract
This paper proposes a novel method to improve locomotion learning for a single quadruped robot using multi-agent deep reinforcement learning (MARL). Many existing methods use single-agent reinforcement learning for an individual robot or MARL for the cooperative task in multi-robot systems. Unlike existing methods, this paper proposes using MARL for the locomotion learning of a single quadruped robot. We develop a learning structure called Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), considering each leg as an agent to explore the action space of the quadruped robot, sharing a global critic, and learning collaboratively. Experimental results indicate that MASQ not only speeds up learning convergence but also enhances robustness in real-world settings, suggesting that applying MASQ to single robots such as quadrupeds could surpass traditional single-robot reinforcement learning approaches. Our study provides insightful guidance on integrating MARL with single-robot locomotion learning.
Index Terms:
Robot control, quadruped robot locomotion, reinforcement learning (RL), multi-agent RL.
@@ I Introduction
Deep reinforcement learning (RL) has made remarkable progress in various robot control learning, such as quadruped robot [1], biped robot [2], and unmanned aerial vehicle [3]. This paper focuses on the quadruped robot control learning.
In the process of applying deep RL control to a single robot, it is prevalent to use single-agent algorithms[4, 5, 6, 7]. However, single-agent algorithms have limits in managing coordination in specific problems. Many existing methods use single-agent RL algorithms for individual robot learning or multi-agent deep reinforcement learning (MARL) for multi-robot systems in cooperative tasks [8, 9, 10]. MARL Cooperation has been widely demonstrated in Game-AI [11, 12] to have significant advantages in multi-agent collaboration.
@@
Unlike existing approaches, we propose using MARL for the locomotion learning of a single quadruped robot to enhance cooperation between its four legs, enabling it to navigate complex terrains and perform intricate tasks. By introducing cooperation MARL, where each leg acts as an agent, the quadruped robot can better coordinate its movements. This collaborative learning structure, termed Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), allows the robot to share experiences. Fig. 1 presents the deployment of the MASQ algorithm on a quadruped robot tested across different terrains, including grass, rocks, flat surfaces, and rubber tracks, demonstrating its performance in diverse environments. Fig.2 shows a sim-to-real comparison of trot gaits in a quadruped robot, highlighting the consistency between simulated and real-world gait patterns.
@@
  • This paper proposes MASQ, a method treats each leg of the single quadruped robot as an agent. The locomotion learning can be modeled and solved as a cooperative MARL problem.
  • Experimental results show that the proposed method enhances performance in executing gaits, outperforms traditional single-agent approaches and improves the efficiency in training and robustness of robotic systems.
@@ II-B MARL for multi-robot control
In multi-agent settings, algorithms like Multi-Agent Proximal Policy Optimization (MAPPO) [32]. Temporally Extended Multi-Agent Reinforcement Learning (TEMP) [33] have demonstrated strong capabilities in addressing multi-agent robotic challenges, such as drone fleet control [34, 35] and autonomous vehicle fleets [36, 37]. Methods like Reinforced Inter-Agent Learning (RIAL) and Differentiable Inter-Agent Learning (DIAL) [38] further enhance collaborative performance by developing communication protocols among agents. This paper proposes modeling single-quadruped robot locomotion learning as a cooperative MARL problem, where each leg is treated as an independent agent, contrasting with previous approaches that treat the robot as a monolithic entity [39, 40, 41, 42] or cooperative groups of robots [43, 44, 45].
@@ III Preliminaries
This paper considers a finite-horizon Markov decision process (MDP) [46], defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,T)$. $\mathcal{S}$ denotes the state space, $\mathcal{A}:=\{a_{0},a_{1},...,a_{|\mathcal{A}|-1}\}$ represents a finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the state transition distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes a reward function, $\gamma\in[0,1)$ denotes a discount factor, and $T$ is a time horizon. At each time step $t$, the policy $\pi$ selects an action $a_{t}\in\mathcal{A}$. After entering the next state by sampling from $\mathcal{P}\left(s_{t+1}\mid s_{t},a_{t}\right)$, the agent receives an immediate reward $r\left(s_{t},a_{t}\right)$. The agent continues to perform actions until it enters a terminal state or $t$ reaches the time horizon $T$. Reinforcement learning (RL) aims to learn a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ for decision-making problems by maximizing discounted rewards. For any policy $\pi$, the state-action value function ($Q$ function) is defined as
@@
Proximal Policy Optimization (PPO) enhances the stability and performance of policy gradient methods by limiting policy updates to prevent destabilizing deviations. The core PPO update rule optimizes a clipped surrogate function:
$L^{\text{CLIP}}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ (2)
@@
$\hat{A}_{t}=R_{t}-V_{\phi}(s_{t})$ (4)
Policy parameters $\theta$ are updated as:
$\theta\leftarrow\theta+\alpha\nabla_{\theta}L^{\text{CLIP}}(\theta)$ (5)
@@
IV Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion Learning
@@
Although many approaches employ deep RL to address single-agent motion control, relying solely on single-agent methods can constrain the complexity and adaptability of learned behaviors in dynamic environments. In this work, we utilize the collaborative potential of multiple agents to improve the learning process for a single robot’s locomotion, enhancing its stability, agility, and adaptability. The proposed method allows multiple virtual agents to independently explore the action space of the quadruped robot, sharing their experiences and learning collaboratively. It enables the robot to benefit from diverse strategies and insights that are difficult to achieve through a single-agent perspective. Specifically, each leg of the quadruped robot is treated as a separate agent within the multi-agent structure, with individual observations and a shared knowledge base, significantly enhancing the cooperation among the robot’s limbs for more effective locomotion.
@@ IV-A MASQ modeling
This paper formulates quadruped robot locomotion learning as a cooperative multi-agent problem, which is described as a partially observable decentralized Markov decision process (decPOMDP) [47]. The decPOMDP is defined by the tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{Z},\mathcal{O},N,\gamma,T)$. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the state transition function, $r$ is the reward function, $\mathcal{Z}$ is the observation space, $\mathcal{O}$ is the observation function, $N$ is the number of agents, $\gamma$ is the discount factor, and $T$ is the time horizon.
At each time step $t$, each agent $n\in\{1,\ldots,N\}$ selects an action $a^{n}_{t}\in\mathcal{A}$, resulting in a joint action $\bm{a_{t}}=\{a^{1}_{t},a^{2}_{t},\ldots,a^{N}_{t}\}$. The environment transitions to a new state $s_{t+1}$ according to $\mathcal{P}(s_{t+1}|s_{t},\bm{a_{t}})$ and provides a shared reward $r(s_{t},\bm{a_{t}})$. Each agent receives an observation $z^{n}_{t}$ from $\mathcal{O}(s_{t},n)$ and maintains an observation-action history $\tau^{n}_{t}$. The goal in MARL is to learn policies $\{\pi^{n}\}_{n=1}^{N}$ that maximize the expected cumulative reward:
@@
MAPPO extends PPO to multi-agent systems, optimizing the following objective function in a multi-agent context:
@@
Each agent updates its policy parameters as follows:
$\theta_{i}\leftarrow\theta_{i}+\alpha\nabla_{\theta_{i}}L^{\text{CLIP}}_{i}(\theta_{i})$ (9)
@@
MAPPO uses centralized training with decentralized execution (CTDE)[48] to handle multi-agent learning challenges. This approach maintains stability in a constantly changing environment. In MAPPO’s implementation, separate neural networks learn a policy ($\pi_{\theta}$) and a value function ($V_{\phi}(s)$) similar to the single-agent version. The value function helps reduce training variability and can take in additional global information.
This paper introduces Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), which applies MARL principles to treat different parts of the robot as independent agents, trained collaboratively using shared rewards. This approach leads to faster learning in simulations and improved robustness in deployment.
IV-B MASQ parameters
Figure 4: Contact force of four legs in different gaits
@@ IV-B MASQ parameters
State Space and Observation: The shared actor network takes a concatenated observation from four agents, each with 35 dimensions, including motor positions $q_{t}\in\mathbb{R}^{3}$, motor speeds $\dot{q_{t}}\in\mathbb{R}^{3}$, previous actions $a_{t-1}\in\mathbb{R}^{3}$ and $a_{t}\in\mathbb{R}^{3}$, a gait sequencing director $d_{t}\in\mathbb{R}^{1}$, projected gravity $g_{t}\in\mathbb{R}^{3}$, command values $v_{t}^{cmd}\in\mathbb{R}^{15}$, body speeds $v_{b}\in\mathbb{R}^{3}$, and a one-hot encoding $e_{t}\in\mathbb{R}^{1}$. The total input for the actor network is $o_{t}^{\text{actor}}\in\mathbb{R}^{140}$ (35x4). The Critic network, using global observations, has a 73-dimensional input, including motor positions $q_{t}\in\mathbb{R}^{12}$, motor speeds $\dot{q_{t}}\in\mathbb{R}^{12}$, previous actions $a_{t-1}\in\mathbb{R}^{12}$ and $a_{t}\in\mathbb{R}^{12}$, gait directors $d_{t}\in\mathbb{R}^{4}$, projected gravity $g_{t}\in\mathbb{R}^{3}$, command values $v_{t}^{cmd}\in\mathbb{R}^{15}$, and body speeds $v_{b}\in\mathbb{R}^{3}$. The output of the Critic network consists of continuous $V$ values $V_{t}\in\mathbb{R}^{4}$, which are used to calculate the advantage function.
@@
Action Space: The output of the actor-network consists of continuous actions $a_{t}\in\mathbb{R}^{12}$, and the system then uses these to calculate the torques for the 12 motors.
@@
Reward Function: The reward functions used in this study are designed to optimize the robot’s performance by encouraging desired behaviors and penalizing undesired ones. Key rewards include: Tracking Linear Velocity, which uses an exponential decay function to minimize velocity error; Linear Velocity Z and Angular Velocity XY, both penalizing unwanted motions to ensure stability; Torques, DOF Velocity, and DOF Acceleration, which promote energy efficiency and smoother movements; and Collision, which penalizes excessive contact forces. Additional rewards focus on action smoothness, accurate jumping, minimizing foot slip and impact velocities, and enhancing locomotion learning stability using Raibert’s heuristic and foot velocity tracking.
Figure 5: Robustness test
TABLE I: Reward settings, corresponding equations, and their values.
Reward Term Equation Value
@@ IV-C Multi-agent actor and global critic networks
In the simulated environment of Isaac Gym, the robot receives observations and rewards to facilitate its learning. The learning process involves dividing the observations into shared and private features, and both the actor and critic networks use the rewards to train the policy.
The actor-network consists of a multi-layer perception (MLP) base and an activation layer that produces actions and their associated log probabilities. Similarly, the critic network uses an MLP base, ending in an output layer that predicts value functions. The actor-network selects actions based on observations, while the critic network assesses the value of these actions to guide the learning process. The actions produced for the agents are processed by an actuator network to simulate real-world conditions, enhancing deployment effectiveness in natural environments. After training, the trained actor-network is deployed onto the robotic dog to perform actions directly. Fig.3 illustrates the entire learning process.
In the context of a quadruped robot, we consider each leg as an individual agent. All four agents share a standard actor-network. Using a shared-parameter network instead of four separate actor networks helps reduce computational load and better fits the nature of the quadruped robot. Unlike typical multi-agent environments, such as Starcraft, where each soldier is an independent agent, the quadruped robot is a single entity with four legs symmetrically positioned around the body’s center. Therefore, a shared-parameter actor-network is more suitable for this scenario.
Because all four legs share the same actor network, we express the policy for each leg as follows:
@@
The quadruped robot has four legs, each with three motors: Hip, Thigh, and Calf joint motors. Each motor’s position, velocity, and torque are observable, so we use these details as the independent observations for each agent. Additionally, to enhance the coordination among the agents, we augment each independent observation with shared observations, including speed and gravity projections calculated from IMU data, temporal director, and agent ID. The temporal director guides the gait sequence of each leg under different movement postures, while the agent ID is necessary for the shared-parameter network. This setup ensures the independence of each agent while improving their cooperative capabilities. The temporal director helps to synchronize the movements of different legs, ensuring smooth gait patterns. It can be defined as:
$T_{i}(t)=\sin\left(2\pi\times F_{i}(t)\right)$ (11)
$F_{i}(t)=G(t)+\phi_{i}+\Delta_{i}$ (12)
where:
  • $F_{i}(t)$ represents the temporal index of the $i$-th leg at time $t$.
  • $G(t)$ is the current phase of the gait cycle, increasing over time and wrapping within $[0,1)$.
  • $\phi_{i}$ is the phase offset associated with the gait for the $i$-th leg.
  • $\Delta_{i}$ is an additional time offset for the $i$-th leg, which is related to the gait.
The Global Critic method uses a centralized value function approach to consider global information, which fits into the CTDE. It relies on a worldwide value function to coordinate individual agents, such as single-agent algorithms like PPO. The Critic network takes in global observations, which consist of the separate observations of the four agents and shared global information.

### v2

@@ Abstract
This paper proposes a novel method to improve locomotion for a single quadruped robot using multi-agent deep reinforcement learning (MARL). Many existing methods use single-agent reinforcement learning for an individual robot or MARL for the cooperative task in multi-robot systems. Unlike existing methods, this paper proposes using MARL for the locomotion of a single quadruped robot. We propose a learning structure called Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), considering each leg as an agent to explore the action space of the quadruped robot, sharing a global critic, and learning cooperatively. Experimental results show that MASQ not only speeds up learning convergence but also enhances robustness in real-world settings, suggesting that applying MASQ to single robots such as quadrupeds could surpass traditional single-robot reinforcement learning approaches. Our study provides insightful guidance on integrating MARL with single-robot locomotion.
@@ I Introduction
Reinforcement learning (RL) has made remarkable progress in various robot control learning [1], such as quadruped robot [2, 3], biped robot [4, 5], and unmanned aerial vehicle [6]. This paper focuses on the quadruped robot control learning.
In the process of applying deep RL to a single robot, it is prevalent to use single-agent algorithms[7, 8, 9, 10]. However, single-agent algorithms may have limitations in managing coordination in specific problems. Many existing methods use single-agent RL algorithms for individual robot learning or multi-agent deep reinforcement learning (MARL) for multi-robot systems in cooperative tasks [11, 12, 13]. Cooperative MARL algorithms have been widely demonstrated in game-AI [14, 15] to have advantages in multi-agent cooperation.
@@
Unlike existing approaches, this paper proposes using MARL for the locomotion of a single quadruped robot to enhance cooperation between its four legs, thereby enabling it to navigate complex terrains and perform intricate tasks. By proposing cooperative MARL, where each leg acts as an agent, the quadruped robot can better coordinate its movements. This collaborative learning structure, termed Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), allows the robot to share experiences. Fig. 1 presents the deployment of the MASQ algorithm on a quadruped robot tested across different terrains, including grass, rocks, flat surfaces, and rubber tracks, demonstrating its performance in diverse environments. Fig. 2 shows a sim-to-real comparison of trot gaits in a quadruped robot, highlighting the consistency between simulated and real-world gait patterns.
@@
  • This paper proposes MASQ, a method that treats each leg of a quadruped robot as an individual agent. The locomotion learning is modeled as a cooperative multi-agent reinforcement learning (MARL) problem and solved using a MARL algorithm.
  • Experimental results show that the proposed method enhances performance in executing gaits, improves training efficiency and robustness, and achieves better final performance, demonstrating the value and impact of the approach.
@@ II-B MARL for Multi-robot Control
In multi-agent settings, algorithms like Multi-Agent Proximal Policy Optimization (MAPPO) [35], Temporally Extended Multi-Agent Reinforcement Learning (TEMP) [36] have demonstrated strong capabilities in addressing multi-agent robotic challenges, such as drone fleet control [37, 38] and autonomous vehicle fleets [39, 40]. Methods like Reinforced Inter-Agent Learning (RIAL) and Differentiable Inter-Agent Learning (DIAL) [41] further enhance collaborative performance by developing communication protocols among agents. This paper proposes modeling single-quadruped robot locomotion as a cooperative MARL problem, where each leg is treated as an independent agent, contrasting with previous approaches that treat the robot as a monolithic entity [42, 43, 44, 45] or cooperative groups of multi-robots [12, 46, 47].
@@ III Preliminaries
This paper considers a finite-horizon Markov decision process (MDP) [48], defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,T)$. $\mathcal{S}$ denotes the state space, $\mathcal{A}:=\{a_{0},a_{1},...,a_{|\mathcal{A}|-1}\}$ represents a finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the staåte transition distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes a reward function, $\gamma\in[0,1)$ denotes a discount factor, and $T$ is a time horizon. At each time step $t$, the policy $\pi$ selects an action $a_{t}\in\mathcal{A}$. After entering the next state by sampling from $\mathcal{P}\left(s_{t+1}\mid s_{t},a_{t}\right)$, the agent receives an immediate reward $r\left(s_{t},a_{t}\right)$. The agent continues to perform actions until it enters a terminal state or $t$ reaches the time horizon $T$. RL aims to learn a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ for decision-making problems by maximizing discounted rewards. For any policy $\pi$, the state-action value function ($Q$ function) is defined as
@@
Proximal Policy Optimization (PPO) [25] enhances the stability and performance of policy gradient methods by limiting policy updates to prevent destabilizing deviations. The core PPO update rule optimizes a clipped surrogate function:
$L^{\textit{CLIP}}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ (2)
@@
$\hat{A}_{t}=Q^{\pi}(s_{t},a_{t})-V_{\psi}(s_{t})$ (4)
where $\psi$ denotes the parameters of value function ($V_{\psi}$) network, $\epsilon$ denotes a coefficient. Policy parameters $\theta$ are updated as:
$\theta\leftarrow\theta+\alpha\nabla_{\theta}L^{\textit{CLIP}}(\theta)$ (5)
@@
IV Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion
@@
In this paper, we use the collaborative potential of multiple agents to improve the learning process for a single robot’s locomotion, resulting in faster training convergence, robustness and better final performance in real-world environments. Specifically, each leg of the quadruped robot is treated as a separate agent within the multi-agent structure, with individual observations and a shared global critic, significantly improving the cooperation among the robot’s limbs for more effective locomotion.
@@ IV-A MASQ Modeling
This paper models a single quadruped robot locomotion as a cooperative multi-agent problem, which is described as a partially observable decentralized Markov decision process (decPOMDP) [49]. The decPOMDP is defined by the tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{Z},\mathcal{O},N,\gamma,T)$. $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the state transition distribution, $r$ is the reward function, $\mathcal{Z}$ is the observation space, $\mathcal{O}$ is the observation function, $N$ is the number of agents, $\gamma$ is the discount factor, and $T$ is the time horizon. At each time step $t$, each agent $n\in\{1,\ldots,N\}$ selects an action $a^{n}_{t}\in\mathcal{A}$, resulting in a joint action $\bm{a_{t}}=\{a^{1}_{t},a^{2}_{t},\ldots,a^{N}_{t}\}$. The environment transitions to a new state $s_{t+1}$ according to $\mathcal{P}(s_{t+1}|s_{t},\bm{a_{t}})$ and provides a shared reward $r(s_{t},\bm{a_{t}})$. Each agent receives an observation $z^{n}_{t}$ from $\mathcal{O}(s_{t},n)$ and maintains an observation-action history $\tau^{n}_{t}$. MARL aims to learn policies $\{\pi^{n}\}_{n=1}^{N}$ that maximize expected cumulative rewards:
@@
This paper proposes Multi-Agent Reinforcement Learning for Single Quadruped Robot Locomotion (MASQ), which applies MARL principles to treat different parts of a single quadruped robot as independent agents, trained collaboratively using shared rewards. Specifically, this paper uses MAPPO [35] to sovle the modeled multi-agent problem. MAPPO optimizes the following objective function in a multi-agent context:
@@
where $i$ denote the $i$-th agent in MARL. Each agent updates its policy parameters as follows:
$\theta_{i}\leftarrow\theta_{i}+\alpha\nabla_{\theta_{i}}L^{\textit{CLIP}}_{i}(\theta_{i})$ (9)
@@
This paper uses centralized training with decentralized execution (CTDE)[50] to handle multi-agent learning challenges. This approach maintains stability in a constantly changing environment. In implementations, separate neural networks learn a policy ($\pi_{\theta}$) and a value function ($V_{\psi}(s)$) similar to the single-agent version. The value function helps reduce training variability and can take in additional global information. This approach leads to better final performance, faster learning speed, and improved robustness in deployment.
IV-B MASQ Settings
@@
State Space and Observation: The shared actor network takes a concatenated observation from four agents, each with 35 dimensions, including motor positions $q_{t}\in\mathbb{R}^{3}$, motor speeds $\dot{q_{t}}\in\mathbb{R}^{3}$, previous actions $a_{t-1}\in\mathbb{R}^{3}$ and $a_{t}\in\mathbb{R}^{3}$, a gait sequencing director $d_{t}\in\mathbb{R}^{1}$, projected gravity $g_{t}\in\mathbb{R}^{3}$, command values $v_{t}^{cmd}\in\mathbb{R}^{15}$, body speeds $v_{b}\in\mathbb{R}^{3}$, and a one-hot encoding $e_{t}\in\mathbb{R}^{1}$. The total input for the actor network is $o_{t}^{\text{actor}}\in\mathbb{R}^{140}$ (35x4). The architecture of the actor network consists of a normalization layer, followed by an MLP layer, a GRU layer, and another normalization layer, with the final output being the joint angle commands for each leg. Additionally, a bias is applied to the output for better precision in control. On the other hand, the critic network uses global observations and has a 73-dimensional input, including motor positions $q_{t}\in\mathbb{R}^{12}$, motor speeds $\dot{q_{t}}\in\mathbb{R}^{12}$, previous actions $a_{t-1}\in\mathbb{R}^{12}$ and $a_{t}\in\mathbb{R}^{12}$, gait directors $d_{t}\in\mathbb{R}^{4}$, projected gravity $g_{t}\in\mathbb{R}^{3}$, command values $v_{t}^{cmd}\in\mathbb{R}^{15}$, and body speeds $v_{b}\in\mathbb{R}^{3}$. The output of the critic network consists of continuous $V$ values $V_{t}\in\mathbb{R}^{4}$, which are used to calculate the advantage function.
@@
Action Space: The output of the actor-network consists of continuous actions $a_{t}\in\mathbb{R}^{12}$, and the system then uses these to calculate the torques for the 12 motors. Details can be found in Section IV-D.
@@
Reward Function: The reward functions in Table I are designed to optimize the robot’s performance by encouraging desired behaviors and penalizing undesired ones. Key rewards include: tracking linear velocity, which uses an exponential decay function to minimize velocity error; linear velocity Z and angular velocity XY, both penalizing unwanted motions to ensure stability; torques, DOF velocity, and DOF acceleration, which promote energy efficiency and smoother movements; and collision, which penalizes excessive contact forces. Additional rewards focus on action smoothness, accurate jumping, minimizing foot slip and impact velocities, and enhancing locomotion stability using Raibert’s heuristic[51] and foot velocity tracking.
TABLE I: Reward function
REWARD SETTINGS, CORRESPONDING EQUATIONS, AND THEIR SCALES
Reward Term Equation Scale
@@ IV-C Multi-agent Actor and Global Critic Networks
In the simulation environment of Isaac Gym [16], the robot receives observations and rewards to facilitate its learning. The learning process involves dividing the observations into shared and private features, and both the actor and critic networks use the rewards to train the policy.
The actor-network consists of a multi-layer perception (MLP) base and an activation layer that produces actions and their associated log probabilities. Similarly, the critic network uses an MLP base, ending in an output layer that predicts value functions. The actor-network outputs actions based on observations, while the critic network assesses the value of these actions to guide the learning process. The actions produced for the agents are processed by an actuator network[52] to simulate real-world conditions, enhancing deployment effectiveness in natural environments. After training, the trained actor-network is deployed onto the robotic dog to perform actions directly. Fig. 3 illustrates the entire learning process.
In the context of a quadruped robot, we consider each leg as an individual agent. All four agents share a standard actor-network. Using a shared-parameter network instead of four separate actor networks helps reduce computational load and better fits the nature of a quadruped robot. Unlike typical multi-agent environments, such as StarCraft [14], where each soldier is an independent agent, the quadruped robot is a single entity with four legs symmetrically positioned around the body’s center. Therefore, a shared-parameter actor-network is more suitable for this scenario.
We express the policy for each leg as follows:
@@
The quadruped robot has four legs, each with three motors: hip, thigh, and calf joint motors. Each motor’s position, velocity, and torque are observable, so we use these details as the independent observations for each agent. Additionally, to enhance the coordination among the agents, we augment each independent observation with shared observations, including speed and gravity projections calculated from inertial measurement unit (IMU) data, temporal director, and agent identifier (ID). The temporal director $T_{i}(t)$ guides the gait sequence of each leg under different movement postures, while the agent ID is necessary for the shared-parameter network. This setup ensures the independence of each agent while improving their cooperative capabilities. The temporal director helps to synchronize the movements of different legs, ensuring smooth gait patterns. It can be defined as:
$T_{i}(t)=\sin\left(2\pi(kt+\Delta_{i})\right)$ (11)
where
  • $k$ is the scaling factor of scaling factor of gait cycle.
  • $\Delta_{i}$ is the phase offset for the $i$-th leg, which determines its relative timing within the gait cycle to ensure coordinated movement.
Figure 4: Robustness test
The Global Critic uses a centralized value function approach to consider global information, which fits into the CTDE. It relies on a global value function to coordinate individual agents, such as single-agent algorithms like PPO. The Critic network takes in global observations, which consist of the separate observations of the four agents and shared global information.